OpenStudy (anonymous):

A collection of nickels, dimes, and quarters totals $8.20. The number of nickels and dimes together is twice the number of quarters. The value of the nickels is one-third the value of the dimes. How many of each kind of coin are there?

OpenStudy (alexwee123):

are you having trouble making the equations?

OpenStudy (anonymous):

yes

OpenStudy (alexwee123):

so n= nickels, d= dimes and q= quarters 0.05n + 0.1d + 0.25q = 8.20 n+d=2q n/3 = d

OpenStudy (pfenn1):

@alexwee123, the Value of the nickels is one third the value of the dimes 0.05n=(1/3)0.10d I think

OpenStudy (alexwee123):

ya i noticed :/

OpenStudy (alexwee123):

forgot about the value part :/

OpenStudy (pfenn1):

\[0.05 n =\left( 0.10d \over 3 \right)\]